# Liouville polarization and the rigidity of their Lagrangian skeleta in dimension $4$

### v1

@@ 4.3. Proof of Theorem 3
In this subsection we first give another proof of Theorem 6 for the ball as domain, that does not use a singular polarization but Biran’s smooth polarization. We then use this result to prove Theorem 3. We start with a lemma that is useful if the polarization of the source is smooth.
Lemma 4.5.
Then there exists an $(\alpha,\alpha^{\prime})$-exact symplectic embedding $\varphi\colon\Sigma\to\overset{\raisebox{-1.29167pt}[0.0pt][-1.29167pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip 6.0mu}{\Sigma^{\prime}}$.
Proof.  A basis of $H_{1}(\Sigma,\mathbb{Z})$ is given by oriented closed curves $\gamma_{1},\dots,\gamma_{2g}$ around the $g$ holes and $b-1$ oriented boundary components $\gamma_{2g+1},\dots,\gamma_{2g+b-1}$, see the red curves in Figure 4.5. Since $g+b-1\leq g^{\prime}$, there exists an embedding $f\colon\Sigma\to\overset{\raisebox{-1.29167pt}[0.0pt][-1.29167pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip 6.0mu}{\Sigma^{\prime}}$ as shown in Figure 4.5.
[…]
Figure 4.5. The embedding $f\colon\Sigma\to\overset{\raisebox{-1.29167pt}[0.0pt][-1.29167pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip 6.0mu}{\Sigma^{\prime}}$ (in the figure it is the inclusion).
Notice that $f_{*}\colon H_{1}(\Sigma)\to H_{1}(\Sigma^{\prime},\partial\Sigma^{\prime})$ is injective, over $\mathbb{Z}$ and hence over $\mathbb{R}$. In our context of oriented surfaces, the de Rham isomorphism gives $H_{1}(\Sigma;\mathbb{R})^{*}=H^{1}(\Sigma)$ and $H_{1}(\Sigma^{\prime},\partial\Sigma^{\prime};\mathbb{R})^{*}=H^{1}_{c}(\overset{\raisebox{-1.29167pt}[0.0pt][-1.29167pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip 6.0mu}{\Sigma^{\prime}})$, where both spaces on the right are the de Rham cohomology groups. Thus, $f^{*}\colon H^{1}_{c}(\overset{\raisebox{-1.29167pt}[0.0pt][-1.29167pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip 6.0mu}{\Sigma^{\prime}})\to H^{1}(\Sigma)$ is surjective.
Since the area of $\Sigma^{\prime}$ is larger than the area of $\Sigma$, Moser’s theorem allows to further assume that $f$ is area-preserving: $f^{*}d\alpha^{\prime}=d\alpha$. Write $\gamma_{i}^{\prime}=f(\gamma_{i})$. Since the classes $[\gamma_{i}]$ are linearly independent in $H_{1}(\Sigma)$ and since $f^{*}$ is surjective, there exists a compactly supported closed $1$-form $\vartheta$ on $\Sigma^{\prime}$ with periods
see e.g. [21, §$\S$ 10.2]. Thus $\phi_{\vartheta}^{1}\circ f$ is a $(\alpha,\alpha^{\prime})$-exact symplectic embedding of $\Sigma$ into $\Sigma^{\prime}$. The proof of the proposition is complete. $\square$
@@
Proof.  The polarization of $\operatorname{\mathbb{C}P}^{2}$ of degree $k$ in Example 2.8 restricts to a tame Liouville polarization $(\Sigma_{k},\lambda_{k})$ of the ball whose skeleton is $\Delta_{k}\cap B^{4}(1)$, so
@@
One readily checks that this Liouville polarization is extendable. Moreover, the area, genus, and number of punctures of $\Sigma_{k}$ are well-known and given by
@@
On the other side, proceeding as in the proof of Proposition 4.1, we take for any $A\in\mathbb{N}$ two points $p_{i}$ in $D(\frac{2}{k})$ and $kA$ points $q_{j}$ in $D(A)$, all with weight $\frac{1}{k}$, and consider the union of the discs $D(\frac{2}{k})\times q_{j}$ and $p_{i}\times D(A)$. Using Lemma 3.5 and resolving all the $2kA$ intersections as in Proposition 3.8, we obtain a smooth Liouville polarization $\Sigma_{k}^{\prime}$ of $D(\frac{2}{k})\times D(A)$ with residue $\frac{1}{k}$. Then
@@
When $k=2$ we can get a better estimate and symplectically embed $B^{4}(1)\backslash\Delta_{2}=B^{4}(1)\backslash\mathbb{R}^{2}$ into the ellipsoid $E(\frac{1}{2},2+\varepsilon)$, like [29]. This is because $\Sigma_{2}$ is a sphere with two punctures, hence has no genus. Thus a cylindrical extension of $\Sigma_{2}$ symplectically embeds into a disc of area $2+\varepsilon$. Since by Example 2.14, $(D(2+\varepsilon),\frac{1}{2})$ is a polarization of $\textnormal{\small SDB}(D(2+\varepsilon),\frac{1}{2})$, Theorem 2.20 shows that $B^{4}(1)\backslash\Delta_{2}$ embeds into $\textnormal{\small SDB}(2+\varepsilon,\frac{1}{2})$. And by [23, Lemma 2.1], $\textnormal{\small SDB}(D(2+\varepsilon),\frac{1}{2})$ is symplectomorphic to $E(2+\varepsilon,\frac{1}{2})$. $\diamond$
@@
Proof.  As above, consider the extendable Liouville polarization $(\Sigma_{mNd},\lambda_{mNd})$ of $B^{4}(1)$ of degree $mNd$, whose skeleton is $\Delta_{mNd}$. We have that ${\mathcal{A}}_{\omega}(\Sigma_{mNd})=mNd$,
@@
In order to produce the required Liouville polarization of the target ellipsoid, consider the symplectic ramified covering
@@
One checks without difficulty that the Liouville polarization $(\Sigma_{m},\lambda_{m})$ of $B^{4}(N(Nd^{2}+d))$ of degree $m$ projects under $\Phi$ to a smooth Liouville polarization of degree $m$ of $E(N,Nd^{2}+d)$ (smoothness requires that $N$ and $Nd^{2}+d$ are coprime, which is equivalent to our assumption that $N$ and $d$ are coprime). This polarization can therefore be seen as a Liouville polarization $(\Sigma^{\prime},\lambda^{\prime})$ of degree $mNd$ of $E(\frac{1}{d},d+\frac{1}{N})$. The polarizing curve $\Sigma^{\prime}$ has area
@@
Proof of Theorem 3. It is shown in [30, §$\S$ 6.2] that for every $\varepsilon>0$ there exists $s_{0}\in\mathbb{R}$ such that for every $s\geq s_{0}$ the ellipsoid $\lambda_{s}E(\frac{1}{s},s)$ of volume $B^{4}(a-\frac{\varepsilon}{2})$ symplectically embeds into $(M,\omega)$. The claim now follows by precomposing this embedding with a scaling of a suitable embedding from Proposition 4.8. $\square$
@@ 5. Lagrangian rigidity
And since $L^{\prime}\subset Z^{4}(a+b)$, its displacement energy $e(L)$ in $\mathbb{C}^{2}$ is $<a+b$, so by Chekanov’s result from [8], $A_{\min}(L^{\prime})\leq e(L)<a+b$. Altogether, $A_{\min}(L)<a+b$, as we wished to prove. $\square$
@@ 6. Legendrian barriers
Let $(M^{3},\xi,\alpha)$ be a contact manifold, $\bigl{(}SM=M\times\mathbb{R}_{>0},d(R\alpha)\bigr{)}$ its symplectization, $\Lambda\subset M$ a Legendrian knot, and $X\subset M$ any subset. Assume that there is no Reeb chord of length $\leq T$ from $\Lambda$ to $\Lambda\cup X$.
@@
where $\Phi^{t}_{\alpha}$ is the Reeb-flow on $(M,\alpha)$. The restriction of $\iota$ to every band $\{p\}\times(0,1]\times[0,\infty)$, $p\in\Lambda$, is symplectic for the the forms $d\tau\wedge dt$ and $d(R\alpha)$ on the domain and the target, respectively. Since there is no Reeb chord of length $\leq T$ from $\Lambda$ to $\Lambda\cup X$, we find $\varepsilon>0$ such that $\iota$ is an embedding of $\Lambda\times(0,1]\times[0,T+\varepsilon)$ into
@@
Take a closed disc $D_{\gamma}$ in $(0,1]\times[0,T+\varepsilon]$ of area $T$ and smooth oriented boundary $\gamma$. Then $\iota(\Lambda\times\gamma)$ is an embedded Lagrangian torus in $V$. The actions of the generators $[\Lambda]$ and $[\gamma]$ of $\pi_{1}(L)$ are $\int_{\Lambda}\alpha=0$ and $\int_{\gamma}\alpha=T$, so ${\mathcal{A}}_{\min}(L,SM)=T$. Moreover, the symplectic disc $\iota(D_{\gamma})$ lies in $V$, has boundary on $L$, and has area $T$. $\square$
@@
$\begin{array}[]{lcl}\bigl{(}S\times\mathbb{R}_{>0},R\;\!\lambda_{S}\bigr{)}&\longrightarrow&\bigl{(}\mathbb{R}^{4}\backslash\{0\},\alpha_{{\operatorname{st}}}\bigr{)}\\[1.99997pt] (s,R)&\longmapsto&R\,s\end{array}$
identifies $S\times(0,1]$ with $U\backslash\{0\}$ and $X\times(0,1]$ with the part in $U\backslash\{0\}$ of the cone over $X$ centered at the origin of $\mathbb{R}^{4}$.
@@
Now take a Legendrian knot $\Lambda$ in $S^{3}(1)$ such that there is no Reeb chord of length $\leq\frac{1}{k}$ from $\Lambda$ to $\Lambda_{k}$. Equivalently, $\Lambda$ is disjoint from $\bigcup_{t\in[0,\frac{1}{k}]}\Phi^{-t}_{R}(\Lambda_{k})$, which is $L$. So $\Lambda$ lies in one of the connected components of $S^{3}(1)\backslash L$ and can be seen as a Legendrian knot in $\bigl{(}D(\frac{1}{k})\times S^{1},\alpha_{\operatorname{st}}+dt\bigr{)}$. But a very classical and elementary argument (see e.g. [1, p. 192]) shows that such a knot has a Reeb chord of length $\leq\frac{1}{k}$ (in fact, of length $<\frac{1}{2k}$).

### v2

@@ 4.4. Proof of Theorem 3
In this subsection we first give another proof of Theorem 6 for the ball as domain, that does not use a singular polarisation but Biran’s smooth polarisation. We then use this result to prove Theorem 3.
@@
Proof.  The polarisation of $\operatorname{\mathbb{C}P}^{2}$ of degree $k$ in Example 2.8 restricts to a tame Liouville polarisation $(\Sigma_{k},\lambda_{k})$ of the ball whose skeleton is $\Delta_{k}\cap B^{4}(1)$, so
@@
One readily checks that this Liouville polarisation is extendable. Moreover, the area, genus, and number of punctures of $\Sigma_{k}$ are well-known and given by
@@
On the other side, proceeding as in the proof of Proposition 4.4, we take for any $A\in\mathbb{N}$ two points $p_{i}$ in $D(\frac{2}{k})$ and $kA$ points $q_{j}$ in $D(A)$, all with weight $\frac{1}{k}$, and consider the union of the discs $D(\frac{2}{k})\times q_{j}$ and $p_{i}\times D(A)$. Using Lemma 3.5 and resolving all the $2kA$ intersections as in Proposition 3.8, we obtain a smooth Liouville polarisation $\Sigma_{k}^{\prime}$ of $D(\frac{2}{k})\times D(A)$ with residue $\frac{1}{k}$. Then
@@
When $k=2$ we can get a better estimate and find a (non-exact) symplectic embedding
$\Phi\colon B^{4}(1)\backslash\Delta_{2}=B^{4}(1)\backslash\mathbb{R}^{2}\,\to\,E(\tfrac{1}{2},2+\varepsilon)$
into the ellipsoid $E(\frac{1}{2},2+\varepsilon)$, like [31]. Indeed, $\Sigma_{2}$ is a sphere with two punctures. Thus a cylindrical extension of $\Sigma_{2}$ symplectically embeds into the disc $D(2+\varepsilon)$ . Since by Example 2.14, $(D(2+\varepsilon),\frac{1}{2})$ is a polarisation of $\textnormal{\small SDB}(D(2+\varepsilon),\frac{1}{2})$, Theorem 2.20 (in its version without exactness assumption and conclusion) shows that $B^{4}(1)\backslash\Delta_{2}$ embeds into $\textnormal{\small SDB}(D(2+\varepsilon),\frac{1}{2})$. And by [24, Lemma 2.1], $\textnormal{\small SDB}(D(2+\varepsilon),\frac{1}{2})$ is symplectomorphic to $E(2+\varepsilon,\frac{1}{2})$. Note that the embedding $\Sigma_{2}\to D(2+\varepsilon)$ is not exact since the boundary circles of $\Sigma_{2}$ in its extension both have area $1$, and so $\Phi$ is not exact. See [31, $\S$ 6.4] for a related discussion. $\diamond$
@@
Proof.  As above, consider the extendable Liouville polarisation $(\Sigma_{mNd},\lambda_{mNd})$ of $B^{4}(1)$ of degree $mNd$, whose skeleton is $\Delta_{mNd}$. We have that ${\mathcal{A}}_{\omega}(\Sigma_{mNd})=mNd$,
@@
In order to produce the required Liouville polarisation of the target ellipsoid, consider the symplectic ramified covering
@@
One checks without difficulty that the Liouville polarisation $(\Sigma_{m},\lambda_{m})$ of $B^{4}(N(Nd^{2}+d))$ of degree $m$ projects under $\Phi$ to a smooth Liouville polarisation of degree $m$ of $E(N,Nd^{2}+d)$ (smoothness requires that $N$ and $Nd^{2}+d$ are coprime, which is equivalent to our assumption that $N$ and $d$ are coprime). This polarisation can therefore be seen as a Liouville polarisation $(\Sigma^{\prime},\lambda^{\prime})$ of degree $mNd$ of $E(\frac{1}{d},d+\frac{1}{N})$. The polarizing curve $\Sigma^{\prime}$ has area
@@
Proof of Theorem 3. It is shown in [32, $\S$ 6.2] that for every $\varepsilon>0$ there exists $s_{0}\in\mathbb{R}$ such that for every $s\geq s_{0}$ the ellipsoid $\lambda_{s}E(\frac{1}{s},s)$ of the same volume as the ball $B^{4}(a-\frac{\varepsilon}{2})$ symplectically embeds into $(M,\omega)$. The claim now follows by precomposing this embedding with a scaling of a suitable embedding from Proposition 4.10. $\square$
@@ 5. Lagrangian rigidity
And since $L^{\prime}\subset Z^{4}(a+b)$, its displacement energy $e(L)$ in $\mathbb{C}^{2}$ is <a+b<a+b, so by Chekanov’s result from [8], $A_{\min}(L^{\prime})\leq e(L)<a+b$. Altogether, $A_{\min}(L)<a+b$, as we wished to prove. $\square$
@@ 6. Legendrian barriers
Let $(M^{3},\xi,\alpha)$ be a contact manifold, $\bigl(SM=M\times\mathbb{R}_{>0},d(R\alpha)\bigr)$ its symplectisation, $\Lambda\subset M$ a Legendrian knot, and $X\subset M$ any subset. Assume that there is no Reeb chord of length $\leq T$ from $\Lambda$ to $\Lambda\cup X$.
@@
where $\Phi^{t}_{\alpha}$ is the Reeb-flow on $(M,\alpha)$. The restriction of $\iota$ to every band $\{p\}\times(0,1]\times[0,\infty)$, $p\in\Lambda$, is symplectic for the forms $d\tau\wedge dt$ and $d(R\alpha)$ on the domain and the target, respectively. Since there is no Reeb chord of length $\leq T$ from $\Lambda$ to $\Lambda\cup X$, we find $\varepsilon>0$ such that $\iota$ is an embedding of $\Lambda\times(0,1]\times[0,T+\varepsilon)$ into
@@
Take a closed disc $D_{\gamma}$ in $(0,1]\times[0,T+\varepsilon]$ of area $T$ and smooth oriented boundary $\gamma$. Then $\iota(\Lambda\times\gamma)$ is an embedded Lagrangian torus in $V$. The actions of the generators $\Lambda]$ and $\gamma]$ of $\pi_{1}(L)$ are $\int_{\Lambda}\alpha=0$ and $\int_{\gamma}\alpha=T$, so ${\mathcal{A}}_{\min}(L,SM)=T$. Moreover, the symplectic disc $\iota(D_{\gamma})$ lies in $V$, has boundary on $L$, and has area $T$. $\square$
@@
[…]
@@
[…]